OpenStudy (anonymous):

Can someone please show me how to divide this radical expression? Teacher didn't explain well enough 10√2 (Yes this is 10√2 over √5) / √5

OpenStudy (saifoo.khan):

\[\frac{10 \sqrt 2}{\sqrt5 } \times \frac{ \sqrt 5}{\sqrt 5}\]

OpenStudy (anonymous):

But that would lead to 10√10 / √10 which √10 and √10 cancel eachother out, leaving just 10.

OpenStudy (anonymous):

and that isn't one of the multiple choice answers

OpenStudy (ash2326):

\[\frac{10 \sqrt {10}}{ {5}}\]

OpenStudy (ajprincess):